For a certain ideal gas R = 25.8 ft-lbf/lbm-0R and k = 1.09 (a) What are the values of cp and cv? (b) What mass of this gas would occupy a volume of 15 cu ft at 75 psia and 800F? (c) If 30   Btu are transferred to this gas at constant volume in (b), what are the resulting temperature and pressure?

a. Cp = ____ Answer in 4 decimal places

Cv = _____ Answer in 4 decimal places

b. m = ______ lb Answer in 2 decimal places

c. T2 = _____

P2 = ______ psia


1
Expert's answer
2021-09-12T18:59:19-0400

"R= 25.8\; \frac{ft \cdot lb}{lb \cdot R} \\

k=1.09 = \frac{c_p}{c_v} \\

v= 15 \;cu \;ft \\

p=75 \;psia"

a. For one mole of ideal gas we know that

"c_p -c_v =R \\

\frac{c_p}{c_v} -1 = \frac{R}{c_v} \\

k -1= \frac{R}{c_v} \\

c_v = \frac{R}{k-1} \\

c_v = \frac{25.8}{1.09-1} = 286.6667 \; BTU/F/lb \\

c_p =c_vk \\

c_p = 286.667 \times 1.09 = 312.4667 \;BTU/F/lb"

b. For ideal gas we know

"pV=mRT \\

m =\frac{pV}{RT} \\

m = \frac{75 \;lb \times 144\;in^2}{in^2 \times ft^2} \times \frac{15 \;ft^3 \times lb \;R}{25.8 ft \;bl \times (80 + 459.67)R} \\

m = \frac{75 \times 144 \times 15}{25.8 \times 539.67} \;lb \\

m = 11.635 \;lb"

c. Q = 30 BTU

"T_1 = 80 °F \\

Q=mc_vΔT \\

Q= mc_v(T_2 -T_1) \\

\frac{Q}{mc_v} +T_1 =T_2 \\

T_2 = \frac{30}{286.67 \times 11.635} +80 \\

T_2 = 80.009 \;°F"

For ideal gas we know

"pV=RT"

at constant volume we know

"\frac{p}{T} = constant \\

\frac{p_1}{T_1} = \frac{p_2}{T_2} \\

p_2 = \frac{p_1}{T_1} \times T_2 \\

p_2 = \frac{75 \times 80.009}{80} \\

p_2 = 75.0084 \;psia"